People of the Inca Empire built a long series of roads and rope bridges in the highlands of Peru. Imagine a couple of brave kids playing on such a bridge. By jumping up and down at a frequency of 0.52 Hz, they set up a standing wave with a node on each end and one in the middie. If the bridge is 47.6 m long, what is the speed of the wave? 49.504 X Your response differs from the correct answer by more than 10%. Double check your calculations. m/s Need Help? Resd
Answer #1

We have

Frequency Fn = n*v/2L

here , n = 2, So,

v = F*L = 0.52*47.6 = 24.75 m/s

speed of the wave is 24.75 m/s.